OpenStudy (anonymous):

The volume of a cube is claimed to be 27 cu.in. correct to within + or -.027 cu.in. Use differentials to estimate the possible error in the measurement of the side of the cube

OpenStudy (agent0smith):

Similar to the last problem, but this time you want to find dw/dV, where w is the width of the side of the cube.

OpenStudy (agent0smith):

dV will be +/- 0.027 cubic in

OpenStudy (anonymous):

and the formula is s^3?

OpenStudy (anonymous):

and s=3 b/c i cubed 27

OpenStudy (anonymous):

@agent0smith this one? :D

OpenStudy (agent0smith):

This one should be easy compared to the other. Yes s=3inches but remember we're only interested in the error in the measurement of the side. So we want ds/dV where ds is the error in the side, dV is the error in volume (+/- 0.027 cubic inches). \[V=s ^{3}\] so s is the cube root of V or V^(1/3) \[s=\sqrt[3]{V} = V ^{\frac{ 1 }{ 3 }}\]

OpenStudy (agent0smith):

So differentiate that to get ds/dV \[\frac{ ds }{ dV }=\frac{ 1 }{ 3 }V ^{\frac{ -2 }{ 3}}\] \[ds=\frac{ 1 }{ 3 }V ^{\frac{ -2 }{ 3}}dV\] And now you can just enter V and dV.

OpenStudy (anonymous):

ahh okay ty <333

OpenStudy (anonymous):

#agent0smith can you help me with 2 more problems? x-x

OpenStudy (agent0smith):

I might but i'm going to bed pretty soon. You should get ds = 0.001 btw, so the error in the side is +/- 0.001 inches.

OpenStudy (anonymous):

aw. okay :(( it's been already past my bed time here xD 5:32 AM here :P

OpenStudy (agent0smith):

haha, 5:32 am here too

OpenStudy (anonymous):

LOL and you're going to bed at this time? xD

OpenStudy (agent0smith):

haha yup. Almost always do.

OpenStudy (anonymous):

eh sometimes I don't get sleep cause of my AP classes xD